OpenStudy (anonymous):

Help please A bowl contains 3 red, 8 blue, and 7 black beads. Margaret randomly selects 3 beads. Find the probability of getting a red, blue, and black marble, in that order. A. dependent; 204/7 B. independent; 7/243 C. independent; 1/68 D. dependent; 7/204

jimthompson5910 (jim_thompson5910):

It sounds like you are not replacing the marbles, so that would make the events dependent. Which means P(red, blue, black) = P(red)*P(blue)*P(black) P(red, blue, black) = P(red)*P(blue, take out 1)*P(black, take out 2) P(red, blue, black) = (3/18)*(8/17)*(7/16) I'll let you finish

OpenStudy (anonymous):

i got 7/? i dont know what the diometor is?

jimthompson5910 (jim_thompson5910):

how did you get 7

OpenStudy (anonymous):

i kept dividing and i got 7

jimthompson5910 (jim_thompson5910):

18*17*16 = 4896

jimthompson5910 (jim_thompson5910):

3*8*7 = 168

OpenStudy (anonymous):

yeah I kept dividing from 168

jimthompson5910 (jim_thompson5910):

168/4896 42/1224 ... divide both parts by 4 21/612 ... divide both parts by 2 7/204 ... divide both parts by 3

jimthompson5910 (jim_thompson5910):

And this is if the beads aren't replaced

OpenStudy (anonymous):

Oh okay